m^3-81m=0 with steps please;)
Found 2 solutions by solver91311, lenny460:
Answer by solver91311(24713)



Factor out an . Factor the difference of two squares that remains. Use the zero product rule to create 3 linear equations in . Solve each to get the expected three roots.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by lenny460(1073)

m^3-81m=0 with steps please;)
m^3 - 81m = 0
Factor out the Greatest Common Factor (GCF) = m
m(m^2 - 81) = 0
Factor a difference between two squares:
m(m + 9)(m - 9) = 0

Subproblem 1
Set the factor 'm' equal to zero and attempt to solve:
m = 0

Subproblem 2
m + 9 = 0
Subtract 9 from each side of the equation.
m + 9 - 9 = 0 - 9
Combine like terms:
9 - 9 = 0
0 - 9 = -9
m = -9

Subproblem 3
m - 9 = 0
Add 9 to each side of the equation.
m - 9 + 9 = 0 + 9
Combine like terms:
-9 + 9 = 0
0 + 9 = 9
m = 9

Solution:
m = (0, -9, 9)



Lennox Obuong
Algebra Student
Email: obuong3@aol.com